# Niel’s Chess: A Quantum Game for Schools and the General Public

### v2

@@ I Introduction
At the heart of the upcoming quantum revolution lies the quantum effect superposition, which means that a physical system is in a state which is a combination of two or more states that are mutually exclusive. For example, according to quantum theory, an atom’s location can be a combination of two or more different locations, somehow as if the atom was in multiple places at the same time. However, when someone tries to observe the atom to see what it looks like being at two or more locations at once, the superposition immediately ’collapses’ and the atom will be found at exactly one of those locations, picked by nature in a truly random fashion; nobody in the universe can predict the outcome with certainty. What is more, superposition leads to the even stranger quantum effect entanglement, where two or more spatially separated physical systems seem to be connected, despite having no apparent physical link between them. It is as if two coins tossed in two far-away cities were destined to (randomly) land on the same side.
@@ III-F Entanglement Move
A conventional piece may join the superposition in which an opponent’s indefinite piece participates, by simultaneously moving to an unoccupied square and staying where it is, provided that at least one of the two indefinite pieces which replace the conventional piece attacks the opponent’s indefinite piece in question. [^8] One piece attacking the other symbolizes that they are ”close”; direct or indirect proximity of two quantum systems is a prerequisite for creating entanglement between them through operations performed in a lab. This is called an ’entanglement move’. Figure 8 (left and right) shows an example where Black’s knight joins the superposition in which White’s indefinite rook on e4 participates. All the participating indefinite pieces must be oriented such that their marks lie on the same diagonal, [^9] It might feel uncomfortable that indefinite pieces must be precisely oriented, but this can be taken as a metaphor for quantum states requiring extreme care and precision, compared to dealing with classical states. and each pair must indicate the same type of superposition (equal or unequal) as the targeted opponent’s pair originally indicated. [^10] Rotating the pieces of an indefinite pair by $45$ degrees does not change the type of superposition, as each mark will face the same upper or lower side of the square as before. Additional conventional pieces may join later, see Figure 9.
Refer to captionRefer to caption
@@ V Results and Discussion
Altogether, the pilot experiment suggests that children as young as 10 years old can learn how to play Niel’s Chess, exposing them to some of the most important principles and thinking patterns of the quantum world at an early age.
As for older children, a recent study shows that high-school students already possess all the necessary background to learn rigorous, comprehensive quantum theory [18]. For this age group, Niel’s Chess can supplement other, curricular materials for quantum technology education. Besides schools, Niel’s Chess can also be an ideal outreach tool, targeting those people in the general public who know the rules of chess but do not wish to delve into rigorous quantum physics.
@@ Acknowledgment
The author thanks Manuel Oriol and Yaiza Aragonés-Soria for helpful feedback and discussions.
@@ Appendix C Variants of the Game
Other variants of Niel’s Chess can be devised, by directing and interpreting the marks on the indefinite pieces in different ways. The game can also be played on smaller boards [21], such as on $4\times 4$, $4\times 5$, $4\times 6$ or $5\times 6$ boards with initial positions shown in Figure 14. From an educational perspective, different quantum phenomena can be illustrated by different rules. For instance, one can exhibit partial collapse by allowing conventional pieces to be on three or more squares at once, or complex amplitudes via fine-grained rotations, or the Pauli-$X$ gate by allowing both pieces of an indefinite pair instance to be rotated by $180$ degrees. The limit is the imagination.
Refer to captionRefer to captionRefer to captionRefer to caption

### v5

@@ I Introduction
At the heart of the upcoming quantum revolution lies the quantum effect superposition, which means that a physical system is in a state which is a combination of two or more states that are mutually exclusive. For example, according to quantum theory, an atom’s location can be a combination of two or more different locations, somehow as if the atom was in multiple places at the same time. However, when someone tries to observe the atom to see what it looks like being at two or more locations at once, the superposition immediately ’collapses’ and the atom will be found at exactly one of those locations, picked by nature in a truly random fashion; nobody in the universe can predict the outcome with certainty. What is more, superposition leads to the even stranger quantum effect entanglement, where two or more distinct, typically spatially separated physical systems seem to be connected, despite having no apparent physical link between them. It is as if two coins tossed in two far-away cities were destined to (randomly) land on the same side.
@@ III-F Entanglement Move
A conventional piece may join the superposition in which an opponent’s indefinite piece participates, by simultaneously moving to an unoccupied square and staying where it is, provided that at least one of the two indefinite pieces which replace the conventional piece attacks the opponent’s indefinite piece in question. [^8] One piece attacking another symbolizes that the latter is ”within reach”; direct or indirect proximity of two quantum systems is a prerequisite for creating entanglement between them through operations performed in a lab. This is called an ’entanglement move’. Figure 8 (left and right) shows an example where Black’s knight joins the superposition in which White’s indefinite rook on e4 participates. All the participating indefinite pieces must be oriented such that their marks lie on the same diagonal, [^9] It might feel uncomfortable that indefinite pieces must be precisely oriented, but this can be taken as a metaphor for quantum states requiring extreme care and precision, compared to dealing with classical states. and each pair must indicate the same type of superposition (equal or unequal) as the targeted opponent’s pair originally indicated. [^10] Rotating the pieces of an indefinite pair by $45$ degrees does not change the type of superposition, as each mark will face the same upper or lower side of the square as before. Additional conventional pieces may join later, see Figure 9.
@@ V Results and Discussion
Altogether, the pilot experiment suggests that children as young as 10 years old can learn how to play Niel’s Chess, exposing them to some of the most important principles and thinking patterns of the quantum world at an early age. As for older children, a recent study shows that high school students already possess all the necessary background to learn rigorous, comprehensive quantum theory [18]. For this latter age group, Niel’s Chess can supplement other, curricular materials for quantum technology education.
Besides schools, Niel’s Chess can also be an ideal outreach tool, targeting those people in the general public who know the rules of chess but do not wish to delve into rigorous quantum physics. Within the community of players, the game can be used as a starting point to discuss quantum principles and clear up possible misconceptions.
@@ Acknowledgment
The author thanks Manuel Oriol and Yaiza Aragonés-Soria for their helpful feedback and discussions, and the anonymous reviewers for their valuable comments and observations.
@@ Appendix C Variants of the Game
Other variants of Niel’s Chess can be devised, by directing and interpreting the marks on the indefinite pieces in different ways. The game can also be played on smaller boards [21], such as on $4\times 4$, $4\times 5$, $4\times 6$ or $5\times 6$ boards with initial positions shown in Figure 14. From an educational perspective, different quantum phenomena can be illustrated by different rules. For instance, one can exhibit partial collapse by allowing conventional pieces to be on three or more squares at once, or complex amplitudes via fine-grained rotations, or the Pauli-$X$ gate by allowing both pieces of an indefinite pair instance to be rotated by $180$ degrees, or even a spin state with real amplitudes via a single rotated mark. The limit is the imagination.